An atom absorbs a photon of wavelength 500 nm and emits another photon of wavelength 700 nm. Find the net energy absorbed by the atom in the process.

Grade:12

1 Answers

Arun
6 years ago
λ base 1 = 500 nm = 500 * 10^–9m, λ base 2 = 700 nm = 700 * 10^–9 m E base 1 – E base 2 = Energy absorbed by the atom in the process. = hc [1/ λ base 1 – 1/ λ base 2] ⇒ 6.63 * 3[1/5 – 1/7] * 10^–19 = 1.136 * 10^–19 J